how many pounds of tea worth $4.60 a pound should be mixed with tea worth $6.50 a pound to get 10 pounds of a mixture worth %5.74 a pound?
Answer by richwmiller(17219)

a+b=10,
4.6*a+6.5*b=5.74*10
a=10-b
4.6*(10-b)+6.5*b=57.4
46.0-4.6b+6.5*b=57.4
1.9*b=11.4
b=6.0
a=10-b
a=4.0 lb. at $4.60
b=6.0 lb. at $6.50
check
4.6*4.0+6.5*6.0=5.74*10
18.4+39.0=57.4
57.4=57.4
ok